# Towards deep physics-informed Kolmogorov–Arnold networks

### v1

@@ 4.3 The Lens of Information Bottleneck Theory
To better understand why the RGA KAN architecture not only achieves superior accuracy compared to baseline cPIKANs but also avoids performance degradation with increasing depth, we turn to Information Bottleneck (IB) theory to analyze the training dynamics of both architecture types. In supervised learning, neural networks aim to reproduce target outputs by progressively forming compressed internal representations of the inputs through their layer activations. According to IB theory, an optimally trained model preserves only the information relevant for reproducing the output while discarding irrelevant input details, effectively forming an “information bottleneck” [65]. This learning process typically unfolds in two distinct phases, namely fitting and diffusion, separated by a phase transition [66, 67, 68]; it is during the diffusion phase that the network develops its generalization capabilities. IB theory has also been applied to analyze the training dynamics of neural networks within the PIML framework [48, 26, 35, 42], and has even been extended to incorporate a third phase within this context, termed diffusion equilibrium [48] or total diffusion [35, 42].
@@
In addition, recent studies have shown that the geometric complexity of the network can provide further insight into its training dynamics [69, 42]. This metric, defined via the discrete Dirichlet energy, is given by
@@
The training dynamics of the RGA KAN architecture, analyzed through the lens of IB theory as studied in the context of PIML [48, 35, 42], reveal a clear progression through all three learning phases. During the initial fitting phase, which spans roughly the first $200$ training iterations, the relative $L^{2}$ error remains nearly constant, while the geometric complexity increases steadily. Simultaneously, the SNR exhibits a brief oscillatory pattern before beginning to decline. In this stage, the network primarily memorizes the training data without significant generalization. A subsequent transition marks the onset of the diffusion phase, characterized by a steadily increasing and fluctuating SNR, a slight decrease in relative $L^{2}$ error and a continued increase in geometric complexity. This phase corresponds to an exploratory stage in which the model identifies more effective learning directions and begins to generalize. Finally, the network enters the diffusion equilibrium phase, during which the SNR reaches a stable, though still oscillatory, plateau, the geometric complexity continues to rise and then also plateaus, and the relative $L^{2}$ error drops sharply, indicating a rapid improvement in generalization and predictive accuracy. Remarkably, this qualitative behavior is consistent across all three network widths. The only systematic difference is the timing of the phase transitions, which occur earlier as the model width increases. This observation aligns with our previous findings: when combined with the proposed initialization, increasing the capacity of the RGA KAN architecture does not lead to divergence but instead improves model accuracy.
In contrast, the behavior of the cPIKAN models differs substantially. While they exhibit a fitting phase similar to that of the RGA KAN, the increase in geometric complexity is far more abrupt and several orders of magnitude larger – an effect previously reported for cPIKANs in [42]. After transitioning to the diffusion phase, these models never reach the diffusion equilibrium phase. The geometric complexity plateaus prematurely, the SNR exhibits strong oscillations without converging to a stable plateau, and the relative $L^{2}$ error stagnates. As a result, the cPIKAN models fail to generalize, explaining the poor performance and divergence observed at large depths for the Allen–Cahn equation.
@@ 5 Experimental Results
Having established that RGA KANs, when initialized with the proposed Glorot-like scheme, remain stable as depth increases and avoid the divergence observed in baseline cPIKANs – both empirically (Section 4.2) and through the lens of IB theory (Section 4.3) – we now turn to a series of forward-PDE benchmarks and ablation studies. In all experiments, we use the exact same hyperparameter and optimization settings as those used in Section 3.3 (number of training iterations, number of collocation points, adaptive training hyperparameters, etc.). These settings are applied uniformly across all architectures considered. The goal of this section is not to perform extensive hyperparameter sweeps to obtain state-of-the-art performance for each PDE individually, but rather to demonstrate that RGA KANs already achieve strong results without any task-specific tuning, using a single, generic configuration.
For the main experiments, we use RGA KANs of width $16$ and depth $12$ (corresponding to $N=6$ RGA blocks). To provide a fair comparison, we also evaluate baseline cPIKANs initialized with the proposed Glorot-like scheme and PirateNets, which represent the current state of the art MLP-based architecture on several PDE benchmarks [23]. To match parameter counts across architectures at the same depth, we adjust widths accordingly: for cPIKANs we use width $18$ and for PirateNets we use width $36$. All experiments are repeated with three random seeds for statistical significance. For RGA KANs, we additionally investigate four variants corresponding to different initializations of the adaptive skip parameters, with $(\alpha,\beta)\in\{0,1\}\times\{0,1\}$. After reporting the benchmark results for each architecture, we identify the $(\alpha,\beta)$ configuration that achieves the lowest error and use it as the reference in the subsequent ablation studies.
The ablation experiments aim to quantify the contribution of each adaptive training component to the overall performance of RGA KANs. To this end, we perform the following sequence: (i) train with RBA alone, disabling all other adaptive techniques; (ii) disable RBA while keeping all other techniques enabled; (iii) disable RBA and RAD while keeping causal training and learning-rate annealing; (iv) disable RBA and causal training while keeping RAD and learning-rate annealing; and finally (v) disable RBA and learning-rate annealing while keeping RAD and causal training. Each configuration is again trained with three different random seeds. The following subsections present the results for each PDE benchmark.
@@ 5.1 Allen–Cahn Equation
The first benchmark considered is the Allen–Cahn equation (see D for details), which has already served as the testbed in previous sections. We begin by examining the effect of different $\left(\alpha,\beta\right)$ initializations on RGA KAN performance. As shown in Table 1, initializing with $\alpha=1$ and $\beta=0$ yields the best results, closely followed by $\alpha=1,\beta=1$. From the three trained instances under this optimal setting, we retain the one achieving the lowest relative $L^{2}$ error ($3.96\times 10^{-4}$) for visualization. Figure 9 compares the model prediction against the reference solution and shows the absolute error field, which remains at most $\mathcal{O}\left(10^{-3}\right)$, demonstrating a good agreement between the two.
@@
Table 1: Results for different RGA KAN $(\alpha,\beta)$ initializations on the Allen–Cahn equation. Reported values are mean $\pm$ standard error over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
@@
Table 2: Performance comparison on the Allen–Cahn equation across different architectures. Reported values are mean $\pm$ standard error over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 1. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
RBA RAD Causal LR Annealing Relative $L^{2}$ Error
$(4.46\pm 0.25)\times 10^{-4}$
$(2.88\pm 2.85)\times 10^{-1}$
$(8.00\pm 2.92)\times 10^{-4}$
$(1.43\pm 0.40)\times 10^{-3}$
$(3.66\pm 1.03)\times 10^{-3}$
$(9.26\pm 1.84)\times 10^{-3}$
Table 3: Ablation study on adaptive training components for the Allen–Cahn equation using the best $(\alpha,\beta)$ initialization from Table 1. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values are mean $\pm$ standard error over three seeds.
Finally, we quantify the contribution of the adaptive training components through ablation studies (Table 3). Using only RBA leads to the largest errors and high variability, suggesting strong sensitivity to weight initialization. Disabling RBA while retaining the other components degrades performance by less than one order of magnitude, indicating that the remaining adaptive techniques are sufficient to preserve stability. Among them, learning-rate annealing has the largest individual impact, followed by causal training, while RAD plays a less dominant role for this PDE.
@@ 5.2 Burgers’ Equation
Table 4: Results for different RGA KAN $(\alpha,\beta)$ initializations on Burgers’ equation. Reported values are mean $\pm$ standard error over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
@@
We next consider Burgers’ equation (see D for details) and follow the same experimental procedure. As summarized in Table 4, the best-performing configuration corresponds to $\alpha=0$ and $\beta=1$, closely followed by $\alpha=1$, $\beta=1$. Although the difference in mean relative $L^{2}$ error between these two configurations is small, the former exhibits a substantially lower standard error, indicating more robustness. We therefore retain $\alpha=0$, $\beta=1$ for the subsequent experiments. Figure 10 shows the predicted solution for the seed with the lowest relative $L^{2}$ error ($2.46\times 10^{-4}$), together with the reference solution and the absolute error field, which retains a maximum value of $\mathcal{O}(10^{-3})$.
@@
Table 5: Performance comparison on Burgers’ equation across different architectures. Reported values are mean $\pm$ standard error over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 4. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
RBA RAD Causal LR Annealing Relative $L^{2}$ Error
@@
Table 6: Ablation study on adaptive training components for Burgers’ equation using the best $(\alpha,\beta)$ initialization from Table 4. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values are mean $\pm$ standard error over three seeds.
Finally, we examine the contribution of each adaptive training technique (Table 6). Enabling only RBA leads to a sharp increase in relative error by nearly two orders of magnitude, though training remains stable. Strikingly, when RBA is disabled but the other adaptive techniques are retained, the error actually drops slightly below the fully adaptive configuration, indicating that RBA may act as a mild hindrance in this specific setting. Among the remaining techniques, RAD resampling has the largest individual impact, followed by causal training and learning-rate annealing. This pattern contrasts with the Allen–Cahn case, highlighting that the relative importance of adaptive components can depend strongly on the PDE at hand.
@@ 5.3 Korteweg–De Vries Equation
Table 7: Results for different RGA KAN $(\alpha,\beta)$ initializations on the Korteweg–De Vries equation. Reported values are mean $\pm$ standard error over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
@@
We then turn to the Korteweg–De Vries equation (see D for details). As summarized in Table 7, the best-performing configuration once again corresponds to $\alpha=0$, $\beta=1$. Using the best-performing seed, with a final relative $L^{2}$ error of $3.21\times 10^{-3}$, we plot the predicted solution alongside the reference and the absolute error field (Figure 11).
Table 8 presents the comparison between architectures. Interestingly, PirateNet diverges in this setting, with a high relative $L^{2}$ error of approximately $7.7\times 10^{-1}$ and very low standard error, indicating consistent failure across seeds. This stands in contrast to the results reported in [23], where PirateNet remained stable at similar depths but larger widths, hinting at a sensitivity to model capacity or hyperparameter choices. The cPIKAN model, while also performing poorly, achieves a lower error than PirateNet, suggesting that the proposed Glorot-like initialization once again helps stabilize training. RGA KANs, by comparison, achieve an error nearly two orders of magnitude lower than both baselines.
@@
Table 8: Performance comparison on the Korteweg–De Vries equation across different architectures. Reported values are mean $\pm$ standard error over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 7. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
RBA RAD Causal LR Annealing Relative $L^{2}$ Error
@@
Table 9: Ablation study on adaptive training components for the Korteweg–De Vries equation using the best $(\alpha,\beta)$ initialization from Table 7. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values are mean $\pm$ standard error over three seeds.
The ablation results are summarized in Table 9. Retaining only RBA leads to significant performance degradation and divergence, as reflected by an increase of nearly two orders of magnitude in error. Conversely, when RBA is disabled and the other adaptive techniques remain active, the model preserves stable training with only a moderate error increase compared to the fully adaptive configuration. Among the remaining methods, RAD again emerges as the most critical, as removing it leads to divergence. The removal of causal training or learning-rate annealing results in final errors on the order of $10^{-2}$, suggesting both play meaningful, but secondary, roles.
@@ 5.4 Sine Gordon Equation
Table 10: Results for different RGA KAN $(\alpha,\beta)$ initializations on the Sine Gordon equation. Reported values are mean $\pm$ standard error over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
@@
We next consider the Sine Gordon equation (see D for details). As shown in Table 10, the configuration with $\alpha=0$ and $\beta=0$ achieves the lowest mean relative $L^{2}$ error, indicating that initializing the network with an effective depth of a single layer is beneficial in this setting. Among the trained instances, the lowest error achieved for this configuration is $2.84\times 10^{-2}$. Figure 12 shows the corresponding prediction, reference solution and absolute error. Notably, the error grows toward the final stages of the temporal domain, suggesting that a more strict tolerance for causal training or additional training iterations could further improve performance.
@@
Table 11: Performance comparison on the Sine Gordon equation across different architectures. Reported values are mean $\pm$ standard error over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 10. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
RBA RAD Causal LR Annealing Relative $L^{2}$ Error
@@
Table 12: Ablation study on adaptive training components for the Sine Gordon equation using the best $(\alpha,\beta)$ initialization from Table 10. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values are mean $\pm$ standard error over three seeds.
Finally, the ablation results are summarized in Table 12. In this case, all ablations lead to relative $L^{2}$ errors of the same order of magnitude as the fully adaptive configuration, with the notable exception of removing learning-rate annealing. In that scenario, the error increases to an average of $1.8\times 10^{-1}$, highlighting the key role of learning-rate annealing for this benchmark.

### v2

@@ 4.3 The Lens of Information Bottleneck Theory
To better understand why the RGA KAN architecture not only achieves superior accuracy compared to baseline cPIKANs but also avoids performance degradation with increasing depth, we turn to Information Bottleneck (IB) theory to analyze the training dynamics of both architecture types. In supervised learning, neural networks aim to reproduce target outputs by progressively forming compressed internal representations of the inputs through their layer activations. According to IB theory, an optimally trained model preserves only the information relevant for reproducing the output while discarding irrelevant input details, effectively forming an “information bottleneck” [IBIntro]. This learning process typically unfolds in two distinct phases, namely fitting and diffusion, separated by a phase transition [IB1, IB2, IB3]; it is during the diffusion phase that the network develops its generalization capabilities. IB theory has also been applied to analyze the training dynamics of neural networks within the PIML framework [IBTheory, RBA, FAIR, KKANs], and has even been extended to incorporate a third phase within this context, termed diffusion equilibrium [IBTheory] or total diffusion [FAIR, KKANs].
@@
In addition, recent studies have shown that the geometric complexity of the network can provide further insight into its training dynamics [GeomCompl, KKANs]. This metric, defined via the discrete Dirichlet energy, is given by
@@
The training dynamics of the RGA KAN architecture, analyzed through the lens of IB theory as studied in the context of PIML [IBTheory, FAIR, KKANs], reveal a clear progression through all three learning phases. During the initial fitting phase, which spans roughly the first $200$ training iterations, the relative $L^{2}$ error remains nearly constant, while the geometric complexity increases steadily. Simultaneously, the SNR exhibits a brief oscillatory pattern before beginning to decline. In this stage, the network primarily memorizes the training data without significant generalization. A subsequent transition marks the onset of the diffusion phase, characterized by a steadily increasing and fluctuating SNR, a slight decrease in relative $L^{2}$ error and a continued increase in geometric complexity. This phase corresponds to an exploratory stage in which the model identifies more effective learning directions and begins to generalize. Finally, the network enters the diffusion equilibrium phase, during which the SNR reaches a stable, though still oscillatory, plateau, the geometric complexity continues to rise and then also plateaus, and the relative $L^{2}$ error drops sharply, indicating a rapid improvement in generalization and predictive accuracy. Remarkably, this qualitative behavior is consistent across all three network widths. The only systematic difference is the timing of the phase transitions, which occur earlier as the model width increases. This observation aligns with our previous findings: when combined with the proposed initialization, increasing the capacity of the RGA KAN architecture does not lead to divergence but instead improves model accuracy.
In contrast, the behavior of the cPIKAN models differs substantially. While they exhibit a fitting phase similar to that of the RGA KAN, the increase in geometric complexity is far more abrupt and several orders of magnitude larger – an effect previously reported for cPIKANs in [KKANs]. After transitioning to the diffusion phase, these models never reach the diffusion equilibrium phase. The geometric complexity plateaus prematurely, the SNR exhibits strong oscillations without converging to a stable plateau, and the relative $L^{2}$ error stagnates. As a result, the cPIKAN models fail to generalize, explaining the poor performance and divergence observed at large depths for the Allen–Cahn equation.
@@ 5 Experimental Results
Having established that RGA KANs, when initialized with the proposed Glorot-like scheme, remain stable as depth increases and avoid the divergence observed in baseline cPIKANs – both empirically (Section 4.2) and through the lens of IB theory (Section 4.3) – we now turn to a series of forward PDE benchmarks and ablation studies. With the exception of the experiments presented in Sections 5.8 and 5.9, we use the exact same hyperparameter and optimization settings as those used in Section 3.3. These settings, the details of which can be found in C.3, are applied uniformly across all architectures considered. The goal of this section is not to perform extensive hyperparameter sweeps to obtain state-of-the-art performance for each PDE individually, but rather to demonstrate that RGA KANs already achieve strong results without any task-specific tuning, using a single, generic configuration.
For the main experiments, we use RGA KANs of width $16$ and depth $12$ (corresponding to $N=6$ RGA blocks). To provide a fair comparison, we also evaluate baseline cPIKANs initialized with the proposed Glorot-like scheme and PirateNets, which represent the current state of the art MLP-based architecture on several PDE benchmarks [PirateNets]. To match parameter counts across architectures at the same depth, we adjust widths accordingly: for cPIKANs we use width $18$ and for PirateNets we use width $36$. All experiments are repeated with three random seeds for statistical significance. For RGA KANs, we additionally investigate four variants corresponding to different initializations of the adaptive skip parameters, with $(\alpha,\beta)\in\{0,1\}\times\{0,1\}$. After reporting the benchmark results for each architecture, we identify the $(\alpha,\beta)$ configuration that achieves the lowest error and use it as the reference in the subsequent ablation studies.
The ablation experiments aim to quantify the contribution of each adaptive training component to the overall performance of RGA KANs. To this end, we perform the following sequence: (i) train with RBA alone, disabling all other adaptive techniques; (ii) disable RBA while keeping all other techniques enabled; (iii) disable RBA and RAD while keeping causal training and LRA; (iv) disable RBA and causal training while keeping RAD and LRA; and finally (v) disable RBA and LRA while keeping RAD and causal training. Each configuration is again trained with three different random seeds. The following subsections present the results for each PDE benchmark.
@@ 5.1 Allen–Cahn Equation
The first benchmark considered is the Allen–Cahn equation (see E.1 for details), which has already served as the testbed in previous sections. We begin by examining the effect of different $\left(\alpha,\beta\right)$ initializations on RGA KAN performance. As shown in Table 1, initializing with $\alpha=1$ and $\beta=0$ yields the best results, closely followed by $\alpha=1,\beta=1$. From the three trained instances under this optimal setting, we retain the one achieving the lowest relative $L^{2}$ error ($3.96\times 10^{-4}$) for visualization. Figure 9 compares the model prediction against the reference solution and shows the absolute error field, which remains at most $\mathcal{O}\left(10^{-3}\right)$.
@@
Table 1: Results for different RGA KAN $(\alpha,\beta)$ initializations on the Allen–Cahn equation. Reported values are mean $\pm$ SEM over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
@@
Table 2: Performance comparison on the Allen–Cahn equation across different architectures. Reported values are mean $\pm$ SEM over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 1. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
RBA RAD Causal LRA Relative $L^{2}$ Error Time / Iter.
$(4.46\pm 0.25)\times 10^{-4}$ 5.51 ms
$(2.88\pm 2.85)\times 10^{-1}$ 5.37 ms
$(8.00\pm 2.92)\times 10^{-4}$ 5.33 ms
$(1.43\pm 0.40)\times 10^{-3}$ 5.30 ms
$(3.66\pm 1.03)\times 10^{-3}$ 5.27 ms
$(9.26\pm 1.84)\times 10^{-3}$ 5.30 ms
Table 3: Ablation study on adaptive training components for the Allen–Cahn equation using the best $(\alpha,\beta)$ initialization from Table 1. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values for relative $L^{2}$ error are mean $\pm$ SEM over three seeds.
Finally, we quantify the contribution of the adaptive training components through ablation studies (Table 3). We include the training time per iteration in this table to demonstrate that the computational overhead of enabling or disabling specific adaptive components is marginal; as shown, the time per iteration remains approximately constant across all configurations. Given this consistent behavior, we omit this column in the ablation tables for the subsequent benchmarks to avoid redundancy. In terms of accuracy, using only RBA leads to the largest errors and high variability, suggesting strong sensitivity to weight initialization. Disabling RBA while retaining the other components degrades performance by less than one order of magnitude, indicating that the remaining adaptive techniques are sufficient to preserve stability. Among them, LRA has the largest individual impact, followed by causal training, while RAD plays a less dominant role for this PDE.
@@ 5.2 Burgers’ Equation
Table 4: Results for different RGA KAN $(\alpha,\beta)$ initializations on Burgers’ equation. Reported values are mean $\pm$ SEM over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
@@
We next consider Burgers’ equation (see E.2 for details) and follow the same experimental procedure. As summarized in Table 4, the best-performing configuration corresponds to $\alpha=0$ and $\beta=1$, closely followed by $\alpha=1$, $\beta=1$. Although the difference in mean relative $L^{2}$ error between these two configurations is small, the former exhibits a substantially lower SEM, indicating more robustness. We therefore retain $\alpha=0$, $\beta=1$ for the subsequent experiments. Figure 10 shows the predicted solution for the seed with the lowest relative $L^{2}$ error ($2.46\times 10^{-4}$), together with the reference solution and the absolute error field.
@@
Table 5: Performance comparison on Burgers’ equation across different architectures. Reported values are mean $\pm$ SEM over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 4. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
RBA RAD Causal LRA Relative $L^{2}$ Error
@@
Table 6: Ablation study on adaptive training components for Burgers’ equation using the best $(\alpha,\beta)$ initialization from Table 4. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values for relative $L^{2}$ error are mean $\pm$ SEM over three seeds.
Finally, we examine the contribution of each adaptive training technique (Table 6). Enabling only RBA leads to a sharp increase in relative error by nearly two orders of magnitude, though training remains stable. Strikingly, when RBA is disabled but the other adaptive techniques are retained, the error actually drops slightly below the fully adaptive configuration, indicating that RBA may act as a mild hindrance in this specific setting. Among the remaining techniques, RAD resampling has the largest individual impact, followed by causal training and LRA. This pattern contrasts with the Allen–Cahn case, highlighting that the relative importance of adaptive components can depend strongly on the PDE at hand.
@@ 5.3 Korteweg–De Vries Equation
Table 7: Results for different RGA KAN $(\alpha,\beta)$ initializations on the Korteweg–De Vries equation. Reported values are mean $\pm$ SEM over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
@@
We then turn to the Korteweg–De Vries equation (see E.3 for details). As summarized in Table 7, the best-performing configuration once again corresponds to $\alpha=0$, $\beta=1$. Using the best-performing seed, with a final relative $L^{2}$ error of $3.21\times 10^{-3}$, we plot the predicted solution alongside the reference and the absolute error field (Figure 11).
Table 8 presents the comparison between architectures. Interestingly, PirateNet diverges in this setting, with a high relative $L^{2}$ error of approximately $7.7\times 10^{-1}$ and very low SEM, indicating consistent failure across seeds. This stands in contrast to the results reported in [PirateNets], where PirateNet remained stable at similar depths but larger widths, hinting at a sensitivity to model capacity or hyperparameter choices. The cPIKAN model, while also performing poorly, achieves a lower error than PirateNet, suggesting that the proposed Glorot-like initialization once again helps stabilize training. RGA KANs, by comparison, achieve an error nearly two orders of magnitude lower than both baselines.
@@
Table 8: Performance comparison on the Korteweg–De Vries equation across different architectures. Reported values are mean $\pm$ SEM over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 7. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
RBA RAD Causal LRA Relative $L^{2}$ Error
@@
Table 9: Ablation study on adaptive training components for the Korteweg–De Vries equation using the best $(\alpha,\beta)$ initialization from Table 7. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values for relative $L^{2}$ error are mean $\pm$ SEM over three seeds.
The ablation results are summarized in Table 9. Retaining only RBA leads to significant performance degradation and divergence, as reflected by an increase of nearly two orders of magnitude in error. Conversely, when RBA is disabled and the other adaptive techniques remain active, the model preserves stable training with only a moderate error increase compared to the fully adaptive configuration. Among the remaining methods, RAD again emerges as the most critical, as removing it leads to divergence. The removal of causal training or LRA results in final errors on the order of $10^{-2}$, suggesting both play meaningful, but secondary, roles.
@@ 5.4 Sine Gordon Equation
Table 10: Results for different RGA KAN $(\alpha,\beta)$ initializations on the Sine Gordon equation. Reported values are mean $\pm$ SEM over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
@@
We next consider the Sine Gordon equation (see E.4 for details). As shown in Table 10, the configuration with $\alpha=0$ and $\beta=0$ achieves the lowest mean relative $L^{2}$ error, indicating that initializing the network with an effective depth of a single layer is beneficial in this setting. Among the trained instances, the lowest error achieved for this configuration is $2.84\times 10^{-2}$. Figure 12 shows the corresponding prediction, reference solution and absolute error. Notably, the error grows toward the final stages of the temporal domain, suggesting that a more strict tolerance for causal training or additional training iterations could further improve performance.
@@
Table 11: Performance comparison on the Sine Gordon equation across different architectures. Reported values are mean $\pm$ SEM over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 10. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
RBA RAD Causal LRA Relative $L^{2}$ Error
@@
Table 12: Ablation study on adaptive training components for the Sine Gordon equation using the best $(\alpha,\beta)$ initialization from Table 10. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values for relative $L^{2}$ error are mean $\pm$ SEM over three seeds.
Finally, the ablation results are summarized in Table 12. In this case, all ablations lead to relative $L^{2}$ errors of the same order of magnitude as the fully adaptive configuration, with the notable exception of removing LRA. In that scenario, the error increases to an average of $1.8\times 10^{-1}$, highlighting the key role of LRA for this benchmark.